Alice and Buster are buying baseball cards. Alice buys 5 packs and 3 individual cards; Buster buys 3 packs and gives his sister 2 cards, he has exactly one half the number of cards as Alice. Hoe many cards are in each pack?
Answer by stanbon(75887)

Alice and Buster are buying baseball cards. Alice buys 5 packs and 3 individual cards; Buster buys 3 packs and gives his sister 2 cards, he has exactly one half the number of cards as Alice. Hoe many cards are in each pack?
-----
Let the number of cards in each pack be "p"
Alice has 5p+3 cards
Buster has 3p-2 cards
----
Equation::
3p-2 = (1/2)(5p+3)
6p-4 = 5p+3
p = 7 cards (# in each pack)
----------------
Cheers,
Stan H.
------------